# Bright ring features and polarization structures in Kerr-Sen black hole images of Sagittarius A* illuminated by radiatively inefficient accretion flows

### v3

@@ Abstract
Using general relativistic radiative transfer (GRRT) simulations, we investigate the bright ring features and polarization structures in images of the Kerr-Sen black hole associated with Sgr A*, as illuminated by 230 GHz thermal synchrotron emission from radiatively inefficient accretion flows (RIAF). Our findings reveal that an increase in the dilaton parameter leads to a shrinking of the bright ring, accompanied by enhancements in both its width and brightness. As the disk thickness grows, the bright ring’s diameter and width both decrease. The brightness enhancement induced by the disk thickness is less prominent than that driven by the dilaton parameter. Comparing with the Event Horizon Telescope (EHT) observational data of SgrA*, we present the allowed ranges of black hole parameters, and find that effects of the disk thickness on the allowed parameter space are more strong than those of the observer’s inclination. Furthermore, we analyze the coefficient $\beta_{2}$ to probe the polarization structure of the black hole images, and reveal that effects of the disk thickness on $\beta_{2}$ are much weaker than those from the dilaton parameter.
@@ I Introduction
The horizon-scale images of M87* and Sgr A* captured by the Event Horizon Telescope (EHT) have opened an unprecedented window for testing gravity in strong-field regimes sgraI; EHTsgraVI; m87I; EventHorizonTelescope:2019ths. The brightness distributions and polarization patterns in black hole images contain a wealth of information about electromagnetic emissions in the vicinity of black holes, offering a powerful means to probe the matter distribution and accretion processes surrounding these celestial objects. This is invaluable for gaining insights into the physics in the strong-field regions and verifying theories of gravity Psaltis_2020; Vagnozzi_2023. Consequently, extensive efforts have been dedicated to investigating the images of various black hole systems Qin:2023nog; zhang2024imageskerrmogblackholes; Chen:2022scf; Wang:2023jop.
@@
The bright ring of emission is one of the most important ingredients in the images captured by EHT, which arises mainly from the strong gravitational lensing of synchrotron emission produced by the accretion flow surrounding the black hole Desire_2025. Actually, this bright ring can be decomposed into finer substructures Johnson:2019ljv, most notably the photon ring which originates from photons completing an infinite number of half-orbits around the black hole. Theoretically, the position of the light ring is entirely determined by the spacetime properties of the black hole and is independent of the properties of the accretion disk Johnson_2020. However, it is difficult to be directly observed in the black hole images because the photon ring is almost overlapped with other subrings and its brightness is obscured by other lower-order images. In addition, the current observational precision also does not allow us to detect the light ring in real astronomical observations and the most likely thing to be observed is the bright ring. Although the bright ring depends not only on black hole parameters but also on the properties of the accretion disk around the black hole, it is necessary to reliably extract the bright ring from the blurred black hole images with finite resolution. This is because the size and brightness of the bright ring can facilitate the capture of key characteristics of the black hole and its accretion disk. With the REx algorithm chael2019simulating, the EHT Collaboration found that the bright ring diameter in the Kerr black hole case remains robust across different imaging methods, while several other features of the ring, such as its width, orientation, asymmetry, and fractional central brightness, depend on the imaging pipeline 2019ApJ...875L...4E; sgraIV; EventHorizonTelescope:2019ggy; achour2025blackholephotonring.
@@
Polarization patterns are another important ingredients in black hole images observed by EHT because because they carry unique information about the magnetic fields and plasma dynamics in the extreme environment surrounding the black hole. The structure of near-horizon linear polarization in EHT images can be quantified by the second azimuthal Fourier mode $\beta_{2}$ of the linear polarization image ehtVII; m87VIII; Palumbo_2020, while the corresponding helicity of the spiral of polarization vectors can be characterized by its phase $\angle\beta_{2}$. The value of $\angle\beta_{2}$ can also be applied to constrain the direction of the near-horizon electromagnetic energy flow of the black hole M87* ehtVII. In addition, $\beta_{2}$ has been shown to not only distinguish whether the magnetic field around the black hole is in the MAD or SANE states, but also further discriminate between different spin values within MAD models Palumbo_2020. Moreover, it has been demonstrated that $\angle\beta_{2}$ shows some sensitivity to the spacetime geometry jiang2023shadowsloopquantumblack. For the black hole Sgr A*, the inclination angle is estimated to be around $30^{\circ}$ 2024ApJ...964L..26E; EHTsgraV; sgraI, which is sufficient to yield rotational symmetry that is accessible to $\beta_{2}$ analysis.
@@
The bright ring features and polarization structures in black hole images also depend on the magnetized accretion flows around the black hole. In general, the full dynamics of accretion flows must resort to general relativistic magnetohydrodynamics (GRMHD) simulations with high precision 2022ApJ...930L..16E; 2019m87V. However, their high computational cost renders a systematic exploration of the vast parameter space facing enormous challenges. A semi-analytic radiatively inefficient accretion flow (RIAF) model is a feasible alternative for describing magnetized accretion flows near black holes Yuan_2003; Pu_2018. For the current EHT targets, M87* and Sgr A*, the surrounding hot plasma is considered to be part of a radiatively inefficient accretion flow, which is typically hot, geometrically thick, and optically thin at an observing frequency of 230 GHz. The corresponding horizon-scale images have been analyzed for Kerr black hole jiang2023shadowsloopquantumblack; Yan:2025mlg; Pu_2018; Pu:2016qak. Therefore, it is highly valuable to investigate the impact of black hole parameters and accretion disk thickness on the bright ring features and polarization structures for other black holes within the RIAF framework.
@@
In this paper, we focus on the Kerr-Sen black hole and study effects of dilaton parameter together accretion disk thickness on bright ring features and polarization structures in the black hole images. The main motivation is that the Kerr-Sen black hole is a solution in string theory, which is widely regarded as a promising candidate for the unified theory of everything. The Kerr-Sen black hole carries charge stemming from the $U(1)$ gauge field in string theory, with this charge accompanied by a dilaton field originating from string compactifications. The black holes in astronomical environments are commonly regarded as being electrically neutral, as oppositely charged particles captured from their surroundings are expected to rapidly screen their charge. However, it is shown that a residual charge may still persist in certain accretion flow scenarios 1978PhRvD..17.1518D. Meanwhile, observations of the supermassive black hole at the center of the Milky Way suggest that Sgr A* may indeed possess a small charge 2018MNRAS.480.4408Z; 2021PhRvL.126d1103B; 2021PhRvD.103j4047K. Thus, it is a worthwhile endeavor to search for signatures of these charges, which are accompanied by dilaton fields, in available astrophysical observations. Studies of the Kerr-Sen black hole have shown that the dilaton parameter does indeed affect the shape of its shadow Wei:2013kza; Guo_2020; Hioki_2008; Uniyal_2017; Tsukamoto:2024asy. However, these studies have not addressed the optical morphology of a Kerr-Sen black hole surrounded by RIAFs. It is therefore important to investigate whether the dilaton parameter influences observable characteristics including the bright ring and polarization pattern in black hole images illuminated by such accretion flows, particularly in the blurred ones.
@@ II MODELING RIAF DYNAMICS around Kerr-Sen black holes
Kerr-Sen black hole is a stationary and axisymmetric solution in Einstein-Maxell-Dilaton-Axion theory, whose metric has the form in the Boyer-Lindquist coordinates Bernard_2016; Ghezelbash_2013; PhysRevLett.74.1276; Sen_1992
@@
Here $M$ and $a$ represent the black hole’s mass and angular momentum, respectively. The dilaton parameter $r_{2}$ is related to the black hole’s electric charge $Q$ and the string tension $\alpha^{\prime}$ by $r_{2}=\frac{\alpha^{\prime}Q^{2}}{8M}$. The electric charge of a Kerr-Sen black hole fundamentally is affected by the axion-photon coupling Banerjee_2021. As the dilaton parameter $r_{2}$ vanishes, the metric (1) reduces to that of Kerr black hole. When the rotation parameter $a$ disappears, the spacetime is reduced to a static spherically symmetric black hole determined exactly by its mass and electric charge, as well as the asymptotic dilaton field YAZADJIEV_1999; PhysRevD.43.3140. It is noted that the Kerr-Sen spacetime Eq. (1) can also be derived through a Newman-Janis transformationNewman:1965tw on the spherically symmetric solution in pure dilaton-coupled gravity YAZADJIEV_1999; PhysRevD.43.3140. The radii of black hole horizons satisfy condition $\Delta=0$ and have the form
@@
In general, black hole images are jointly determined by both the black hole parameters and the surrounding plasma Younsi_2023; _zel_2022. For the current observed targets of the EHT including M87* and Sgr A*, their surrounding hot plasmas are considered to be part of a RIAF. Here, we adopt a semi-analytical optically thin RIAF model Yuan_2003; Pu_2018 around the Kerr-Sen black hole to perform general relativistic ray-tracing numerical simulations Noble:2007zx; moscibrodzka2017ipolesemianalyticscheme. In this semi-analytical RIAF model, the electron number density and temperature are assumed to follow power-law profiles with radius,
@@
where $\delta$ and $\gamma$ are the power-law indices for the electron density and the electron temperature, respectively. The temperature index $\gamma$ is observationally constrained by the spectral energy distribution Broderick:2010kx and VLBI brightness temperature measurements Yfantis:2024eab. The parameter $H$ describes the geometric thickness of the disk. Here, we adopt $\delta=1.1$ and $\gamma=0.84$, which are consistent with these observational constraints and previous studies of RIAFs Pu_2018. Since the magnetically arrested disk (MAD) model is favored by the current EHT observations EHTsgraV, here a MAD-like configuration Chen_2022; jiang2023shadowsloopquantumblack is adopted in our simulations.
To model the four-velocity of the accretion flow, one can interpolate between Keplerian orbital motion and geodesic free fall as described in Pu_2018; Pu:2016qak
@@
where $(u^{r},\Omega)$ and $(u^{r}_{\rm ff},\Omega_{\rm ff})$ denote the radial component of the four-velocity $u^{r}$ and the angular velocity $\Omega=u^{\phi}/u^{t}$ of flows moving along Keplerian and free-fall orbits, respectively. The coefficient $\kappa$ is a regulating parameter. Specifically, when $(\kappa_{\rm ff},\kappa_{\rm K})=(0,1)$ and $(\kappa_{\rm ff},\kappa_{\rm K})=(1,0)$, the motion of flows corresponds to pure Keplerian motion and free fall motion, respectively. Here, we set to $(\kappa_{\rm ff},\kappa_{\rm K})=(0.5,0.5)$, which means that the flows are in the sub-Keplerian state. Moreover, here the synchrotron emissivity in the disk is modeled by a relativistic thermal (Maxwell-Jüttner) electron distribution because non-thermal emission is less dominant for the 230 GHz band. In the polarized radiative transfer process, the related coefficients, including emission, absorption, and Faraday effects, are computed using the fitting formulas in Pandya:2016qfh. To ensure that the observed flux of thermal synchrotron radiation matches the measured value of Sgr A*, the normalized electron number density $n_{e,0}$ and the electron temperature $T_{e,0}$ are set to $n_{e,0}\approx 10^{7}\rm cm^{-3}$ and $T_{e,0}\approx 10^{11}\rm K$, respectively Pu_2018. The detail fiducial values are listed in Table 1.

### v4

@@ Abstract
Using general relativistic radiative transfer (GRRT) simulations, we investigate the bright ring features and polarization structures in images of the Kerr-Sen black hole associated with Sgr A*, as illuminated by 230 GHz thermal synchrotron emission from radiatively inefficient accretion flows (RIAF). Our findings reveal that an increase in the dilaton parameter leads to a shrinking of the bright ring, accompanied by enhancements in both its width and brightness. As the disk thickness grows, the bright ring’s diameter and width both decrease. The brightness enhancement induced by the disk thickness is less prominent than that driven by the dilaton parameter. Comparing with the Event Horizon Telescope (EHT) observational data of SgrA*, we present the allowed ranges of black hole parameters, and find that effects of the disk thickness on the allowed parameter space are stronger than those of the observer’s inclination. Furthermore, we analyze the coefficient $\beta_{2}$, associated with the two-fold rotational symmetry of the electric vector position angles (EVPA), to probe the polarization structure of the black hole images, and reveal that effects of the disk thickness on $\beta_{2}$ are much weaker than those from the dilaton parameter.
@@ I Introduction
The horizon-scale images of M87* and Sgr A* captured by the Event Horizon Telescope (EHT) have opened an unprecedented window for testing gravity in strong-field regimes Collaboration et al. (2022b, a); Akiyama et al. (2019b); Akiyama and others (2019a). The brightness distributions and polarization patterns in black hole images contain a wealth of information about electromagnetic emissions in the vicinity of black holes, offering a powerful means to probe the matter distribution and accretion processes surrounding these celestial objects. This is invaluable for gaining insights into the physics in the strong-field regions and verifying theories of gravity Psaltis et al. (2020); Vagnozzi et al. (2023). Consequently, extensive efforts have been dedicated to investigating the images of various black hole systems Qin et al. (2023); Zhang et al. (2024); Chen et al. (2023); Wang et al. (2024).
@@
The bright ring of emission is one of the most important ingredients in the images captured by EHT, which arises mainly from the strong gravitational lensing of synchrotron emission produced by the accretion flow surrounding the black hole Desire et al. (2025). Actually, this bright ring can be decomposed into finer substructures Johnson and others (2020), most notably the photon ring which originates from photons completing an infinite number of half-orbits around the black hole. Theoretically, the position of the light ring is entirely determined by the spacetime properties of the black hole and is independent of the properties of the accretion disk Johnson et al. (2020). However, it is difficult to be directly observed in the black hole images because the photon ring is almost overlapped with other subrings and its brightness is obscured by other lower-order images. In addition, the current observational precision also does not allow us to detect the light ring in real astronomical observations and the most likely thing to be observed is the bright ring. Because the bright ring depends not only on black hole parameters but also on the properties of the accretion disk around the black hole, it is necessary to reliably extract the bright ring from the blurred black hole images with finite resolution. This is because the size and brightness of the bright ring can facilitate the capture of key characteristics of the black hole and its accretion disk. With the REx algorithm Chael (2019), the EHT Collaboration found that the bright ring diameter in the Kerr black hole case remains robust across different imaging methods, while several other features of the ring, such as its width, orientation, asymmetry, and fractional central brightness, depend on the imaging pipeline Event Horizon Telescope Collaboration et al. (2019); Collaboration et al. (2022c); Akiyama and others (2019b); Achour et al. (2025).
@@
Polarization patterns are another important ingredients in black hole images observed by EHT because because they carry unique information about the magnetic fields and plasma dynamics in the extreme environment surrounding the black hole. The structure of near-horizon linear polarization in EHT images can be quantified by the second azimuthal Fourier mode $\beta_{2}$ of the linear polarization image Akiyama et al. (2021a, b); Palumbo et al. (2020), while the corresponding helicity of the spiral of polarization vectors can be characterized by its phase $\angle\beta_{2}$. The value of $\angle\beta_{2}$ can also be applied to constrain the direction of the near-horizon electromagnetic energy flow of the black hole M87* Akiyama et al. (2021a). In addition, $\beta_{2}$ has been shown to not only distinguish whether the magnetic field around the black hole is in the magnetically arrested disk (MAD) or standard and normal evolution (SANE) states, but also further discriminate between different spin values within MAD models Palumbo et al. (2020). Moreover, it has been demonstrated that $\angle\beta_{2}$ shows some sensitivity to the spacetime geometry Jiang et al. (2023). For the black hole Sgr A*, the inclination angle is estimated to be around $30^{\circ}$ Event Horizon Telescope Collaboration et al. (2024); Collaboration et al. (2022d, b). This low inclination is sufficient to yield rotational symmetry that is accessible to $\beta_{2}$ analysis.
@@
The bright ring features and polarization structures in black hole images also depend on the magnetized accretion flows around the black hole. In general, the full dynamics of accretion flows must resort to general relativistic magnetohydrodynamics (GRMHD) simulations with high precision Event Horizon Telescope Collaboration et al. (2022); Akiyama et al. (2019a). However, their high computational cost renders a systematic exploration of the vast parameter space facing enormous challenges. A semi-analytic radiatively inefficient accretion flow (RIAF) model is a feasible alternative for describing magnetized accretion flows near black holes Yuan et al. (2003); Pu and Broderick (2018). For the current EHT targets, M87* and Sgr A*, the surrounding hot plasma is considered to be part of a radiatively inefficient accretion flow, which is typically hot, geometrically thick, and optically thin at an observing frequency of 230 GHz. The corresponding horizon-scale images have been analyzed for Kerr black hole Jiang et al. (2023); Yan et al. (2025); Pu and Broderick (2018); Pu et al. (2016). Therefore, it is highly valuable to investigate the impact of black hole parameters and accretion disk thickness on the bright ring features and polarization structures for other black holes within the RIAF framework.
@@
In this paper, we focus on the Kerr-Sen black hole and study effects of dilaton parameter together with accretion disk thickness on bright ring features and polarization structures in the black hole images. The motivation is based on the following aspects: Firstly, the Kerr-Sen black hole is a solution in string theory, which is widely regarded as a promising candidate for the unified theory of everything. Secondly, the Kerr-Sen metric exhibits substantial differences in its intrinsic geometric properties from the Kerr-Newman solution in general relativity, despite their strong resemblance. Furthermore, investigating this background’s impact on black hole images not only provides a possibility for testing string theory but also offers a scope to verify the Kerr hypothesis, especially given the future development of high-resolution imaging via very long baseline interferometry (VLBI). The Kerr-Sen black hole carries charge stemming from the $U(1)$ gauge field in string theory, with this charge accompanied by a dilaton field originating from string compactifications. Unlike the standard electromagnetic charge, the $U(1)$ charge of Kerr-Sen spacetime essentially originates from the axion-photon coupling rather than infalling charged particles. When the charge vanishes, one can find the field corresponding to both the axion and the dilaton disappear, a conclusion that can also be obtained from Eq.(3). The black holes in astronomical environments are commonly regarded as being electrically neutral, as oppositely charged particles captured from their surroundings are expected to rapidly screen their charge. However, it is shown that a residual charge may still persist in certain accretion flow scenarios Damour et al. (1978). Meanwhile, observations of the supermassive black hole at the center of the Milky Way suggest that Sgr A* Zajaček et al. (2018); Bozzola and Paschalidis (2021); Kocherlakota et al. (2021) may indeed possess a small charge, with an upper observational limit $q<3\times 10^{8}C$ Zajaček et al. (2018). Thus, it is a worthwhile endeavor to search for signatures of these charges, which are accompanied by dilaton fields, in available astrophysical observations. Studies of the Kerr-Sen black hole have shown that the dilaton parameter does indeed affect the shape of its shadow Wei and Liu (2013); Guo et al. (2020); Hioki and Miyamoto (2008); Uniyal et al. (2017); Tsukamoto and Kase (2024); Yagi et al. (2024); Ben Achour et al. (2025). However, these studies have not addressed the optical morphology of a Kerr-Sen black hole surrounded by RIAFs. It is therefore important to investigate whether the dilaton parameter influences observable characteristics including the bright ring and polarization pattern in black hole images illuminated by such accretion flows, particularly in the blurred ones.
@@ II MODELING RIAF DYNAMICS around Kerr-Sen black holes
The Kerr-Sen black hole is a stationary and axisymmetric solution in Einstein-Maxell-Dilaton-Axion theory, whose metric has the form in the Boyer-Lindquist coordinates Bernard (2016); Ghezelbash and Siahaan (2013); García et al. (1995); Sen (1992)
@@
In the Kerr-Sen solution, the axion, dilaton and $U(1)$ gauge fields admit the explicit forms Sen (1992)
$\xi=\frac{q^{2}}{M}\frac{a\cos\theta}{r^{2}+a^{2}\cos^{2}\theta},\quad\quad e^{2\chi}=\frac{r^{2}+a^{2}\cos^{2}\theta}{r(r+r_{2})+a^{2}\cos^{2}\theta},\quad\quad A=\frac{qr}{\tilde{\Sigma}}\bigg(-dt+a\mathrm{sin}^{2}\theta d\phi\bigg).$ (3)
Here $M$ and $a$ represent the black hole’s mass and angular momentum, respectively. The dilaton parameter $r_{2}$ is related to the black hole’s electric charge $q$ and the inverse string tension $\alpha^{\prime}$ by $r_{2}={\frac{\alpha^{\prime}}{8}}\frac{q^{2}}{M}$. Theoretically, string theory predicts $r_{2}$ to be of the order of the square of the Planck length $l^{2}_{p}$ since the inverse string tension $\alpha^{\prime}\simeq l^{2}_{p}$ Polchinski (1998). Here, we follow the standard phenomenological approach in black hole imaging studies Mizuno et al. (2018); Röder et al. (2023); Dastan et al. (2016) and treat $r_{2}$ as a free parameter to investigate its potential effects deviating from general relativity. The origin of the electric charge in the Kerr-Sen solution is not from the falling charged particles, but from the axion-photon coupling Banerjee et al. (2021); Tripathi et al. (2021). As the dilaton parameter $r_{2}$ vanishes, the metric (1) reduces to the Kerr black hole metric instead of the Kerr-Newman one, implying that the charge of the Kerr-Sen black hole is essentially different from that of pure electromagnetic origin. When the rotation parameter $a$ disappears, the spacetime is reduced to a static spherically symmetric black hole determined exactly by its mass and electric charge, as well as the asymptotic dilaton field YAZADJIEV (1999); Garfinkle et al. (1991). It is noted that the Kerr-Sen spacetime Eq. (1) can also be derived through a Newman-Janis transformationNewman and Janis (1965) on the spherically symmetric solution in pure dilaton-coupled gravity YAZADJIEV (1999); Garfinkle et al. (1991). The radii of black hole horizons satisfy condition $\Delta=0$ and have the form
@@
Black hole images are jointly determined by both the black hole parameters and the surrounding plasma Younsi et al. (2023); Özel et al. (2022). In general, the axion, dilaton and electromagnetic fields of Kerr-Sen black hole interact with the disk. However, the high complexity of such interactions makes a comprehensive analysis challenging, even with high-performance computing clusters at one’s disposal. For simplicity, in this work, we only consider the influence of the spacetime metric on the accretion disk as in Mizuno et al. (2018); Fromm et al. (2021); Röder et al. (2023). For the current observed targets of the EHT including M87* and Sgr A*, their surrounding hot plasmas are considered to be part of a RIAF. Here, we adopt a semi-analytical optically thin RIAF model Yuan et al. (2003); Pu and Broderick (2018) around the Kerr-Sen black hole to perform general relativistic ray-tracing numerical simulations Noble et al. (2007); Moscibrodzka and Gammie (2017). In this semi-analytical RIAF model, the electron number density and temperature are assumed to follow power-law profiles with radius,
@@
where $\delta$ and $\gamma$ are the power-law indices for the electron density and the electron temperature, respectively. The temperature index $\gamma$ is observationally constrained by the spectral energy distribution Broderick et al. (2011) and VLBI brightness temperature measurements Yfantis et al. (2024). The parameter $H$ describes the geometric thickness of the disk. Here, we adopt $\delta=1.1$ and $\gamma=0.84$, which are consistent with these observational constraints and previous studies of RIAFs Pu and Broderick (2018). Since the magnetically arrested disk (MAD) model is favored by the current EHT observations Collaboration et al. (2022d), here a MAD-like configuration Chen et al. (2022); Jiang et al. (2023) is adopted in our simulations.
To model the four-velocity of the accretion flow, one can interpolate between Keplerian orbital motion and geodesic free fall as described in Pu and Broderick (2018); Pu et al. (2016)
@@
where $(u^{r},\Omega)$ and $(u^{r}_{\rm ff},\Omega_{\rm ff})$ denote the radial component of the four-velocity $u^{r}$ and the angular velocity $\Omega=u^{\phi}/u^{t}$ of flows moving along Keplerian and free-fall orbits, respectively. The coefficient $\kappa$ is a regulating parameter. Specifically, when $(\kappa_{\rm ff},\kappa_{\rm K})=(0,1)$ and $(\kappa_{\rm ff},\kappa_{\rm K})=(1,0)$, the motion of flows corresponds to pure Keplerian motion and free fall motion, respectively. Here, we set to $(\kappa_{\rm ff},\kappa_{\rm K})=(0.5,0.5)$, which means that the flows are in the sub-Keplerian state. Moreover, here the synchrotron emissivity in the disk is modeled by a relativistic thermal (Maxwell-Jüttner) electron distribution because non-thermal emission is less dominant for the 230 GHz band. In the polarized radiative transfer process, the related coefficients, including emission, absorption, and Faraday effects, are computed using the fitting formulas in Pandya et al. (2016). To ensure that the observed flux of thermal synchrotron radiation matches the measured value of Sgr A*, the normalized electron number density $n_{e,0}$ and the electron temperature $T_{e,0}$ are set to $n_{e,0}\approx 10^{7}\rm cm^{-3}$ and $T_{e,0}\approx 10^{11}\rm K$, respectively Pu and Broderick (2018). The detail fiducial values are listed in Table 1.
@@
In Fig. 1, we present some black hole images and the corresponding blurred counterparts for Kerr-Sen black holes, based on this semi-analytical optically thin RIAF model.